Are the following lines parallel, perpendicular, or neither?
L1 through (9,2) and (2,-8)
L2 through (-3,5) and (5,-1)
-8/2 - 2/9= -10/-7
-1/5 - 5/-3= -6/8
Is this problem neither?
Thank You!
Thank you!

Answer by Nate(3500)

Find the slope of each line:
(9,2) and (2,-8)
M = (y - y)/(x - x) = (9 - 2)/(2 - (-8)) = 7/10
The slope of L1 is (7/10)
(-3,5) and (5,-1)
M = (y - y)/(x - x) = ((-3) - 5)/(5 - (-1)) = (-8)/(6) = (-4/3)
The slope of the lines are not parallel or perpendicular, so the answer is neither.